Find the midpoint of AB if A (-3,8) and B (-7,-6)


1
Expert's answer
2021-08-27T12:44:28-0400

x points are -3 and -7.

( -3+(-7))/2 = -10/2 =-5

y points are 8 and -6.

(8+(-6))/2 = 2/2 =1


The midpoint of AB if A (-3,8) and B (-7,-6) is (-5,1).